Two trains leave the station at the same time, but headed in different directions. The first train at 10 mph faster than the second train. After six hours they are 720 mi apart. What are the speeds of the trains?
Answer by richwmiller(17219)

6*x+6(x+10)=720
12x+60=720
12x=660
2x=110
x=55
x+10=65